OpenStudy (anonymous):

Solve: -4(2x - 1) = -6(x + 2) - 2

Directrix (directrix):

We will need the Distributive Property.

OpenStudy (anonymous):

Thats not one of the options @Directrix /:

Directrix (directrix):

I will do this part: -4(2x - 1) OR this part: -6(x + 2) - 2 and then you the other. Choose which part you'd like me to multiply out. @Nurialozza96

Directrix (directrix):

The Distributive Property is a tool used to assist in determing the answer. It is not the answer.

OpenStudy (anonymous):

@Directrix it has options like ... " A. add 6 to both sides" those kind of steps

Directrix (directrix):

@Nurialozza96 Please post the instructions so that I can understand what we are supposed to be doing. After I read them, then you can delete them.

OpenStudy (anonymous):

@Directrix In order to solve the equation -3x - 6 = 5, the first step is to: add 6 to both sides divide by 5 on both sides subtract 6 from both sides multiply by 5 on both sides

OpenStudy (owlcoffee):

Close, but no. in order to solve \[-3x-6=5\] You first need to: add 6 to both sides: \[+6-3x-6=5+6\] The 6s will simplify. Next you should divide each side by -3 in order to form a 1: \[\frac{ -3x }{ -3 }=\frac{ 5-6 }{ -3 }\] The (-3)s will simplify and we get: \[x=\frac{ 5-6 }{ -3 }\] So we can conclude by addition and substraction that: \[x=\frac{ 1 }{ 3 }\]

OpenStudy (anonymous):

Solve: -4(2x - 1) = -6(x + 2) - 2 -8x+4=-6x -12 -2 -2x=-14-4 -2x=-18 x=-18/-2 x=9

Directrix (directrix):

@Owlcoffee Where did you get -3x - 6 = 5?

Directrix (directrix):

@Nurialozza96 The options you posted do not go with the problem you posted. -4(2x - 1) = -6(x + 2) - 2 is the posted problem. The options talk about: In order to solve the equation -3x - 6 = 5 So, I don't know if two different problems have been posted or what. I guess I'm confused.

OpenStudy (owlcoffee):

I read it one of the comments so I thought it was it.